OpenStudy (anonymous):

Solve the following system of equations: y = 3x – 2 6x – y = 2

OpenStudy (anonymous):

well you know that y = 3x - 2. So sub 3x - 2 in for y in the other equation and solve for x.

OpenStudy (anonymous):

once subbed in, it looks like this... 6x - y = 2 6x - (3x - 2) = 2 -- distribute through the parenthesis 6x - 3x + 2 = 2 3x = 0 x = 0 6x - y = 2 6(0) - y = 2 -y = 2 y = -2 check... y = 3x - 2 -2 = 3(0) - 2 -2 = -2 (correct) x = 0 and y = -2

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

thank you i understand it know

OpenStudy (anonymous):

good to hear :)